# Robust aluminum nitride passivation of silicon carbide with near-surface spin defects

### v1

@@ Abstract
Silicon carbide (SiC) hosts a number of point defects that are being explored as single-photon emitters for quantum applications. Unfortunately, these quantum emitters lose their photostability when placed in proximity to the surface of the host semiconductor. In principle, a uniform passivation of the surface’s dangling bonds by simple adsorbates, such as hydrogen or mixed hydrogen/hydroxyl groups, should remove detrimental surface effects. However, the usefulness of atomic and molecular passivation schemes is limited by their lack of long-term chemical and/or thermal stability. In this first principles work, we use aluminum nitride (AlN) to passivate SiC surfaces in a core-shell nanowire model. By using a negatively charged silicon vacancy in SiC as the proof-of-principle quantum emitter, we show that AlN-passivation is effective in removing SiC surface states from the band gap and in restoring the defect’s optical properties. We also report the existence of novel silicon vacancy-based defects at the SiC-AlN interface.
@@ Introduction
Quantum bits (qubits) and single photon emitters are essential building blocks for quantum information science applications. Wide band gap semiconductors with fluorescent, spin-active point defects can furnish both qubits and single photon emitters for quantum technologies. This is particularly true for different bright defects in silicon carbide (SiC), which not only emit single photons in the favorable near-telecom range Christle et al. (2015); Fuchs et al. (2015), but they also possess addressable spins Falk et al. (2013); Widmann et al. (2015); Seo et al. (2016); Nagy et al. (2019) with long coherence times Seo et al. (2016); Christle et al. (2015); Carter et al. (2015); Nagy et al. (2019); Widmann et al. (2015). The host material itself offers additional advantages, such as low cost, commercial availability, and ease of nanofabrication, all of which can enable scalable SiC-based quantum technologies.
@@
In order to enhance the radiative signal from defects for quantum computing and/or enhance their sensitivity for quantum sensing applications, different strategies are used, such as shallow implantation and/or nanofabrication of optical cavities in the host materials Gadalla et al. (2021); Majety et al. (2021); Crook et al. (2020); Bracher et al. (2017); Nagy et al. (2018); Lukin et al. (2020); Radulaski et al. (2017). A consequence of these strategies is the near-surface placement of the defect, which inadvertently modulates its properties. This has been reported in experiments Lukin et al. (2020); Yuan et al. (2020); Sangtawesin et al. (2019); Vásquez et al. (2020); Neethirajan et al. (2023) and explored in a few theoretical works Kaviani et al. (2014); Li et al. (2019); Löfgren et al. (2019); Joshi and Dev (2022); Sayou Ngomsi et al. (2024). In particular, two recent theoretical works Joshi and Dev (2022); Sayou Ngomsi et al. (2024) studied properties of the near-surface, negatively-charged silicon monovacancy () in a 2H-SiC nanowire (NW). In their first principles study of defects in an unpassivated NW, Joshi and Dev Joshi and Dev (2022) demonstrated that the changes to the frequency of quantum emission from and the loss of the defect’s photostability result from the spatial and energetic proximity of the surface states to the defect states. Sayou Ngomsi et al. Sayou Ngomsi et al. (2024) further showed that the aforementioned detrimental surface effects can be remedied by passivation of the SiC surfaces with simple adsorbates, such as hydrogen or mixed hydrogen/hydroxyl groups. The resulting removal of the SiC surface states from the band gap leads to a near-perfect restoration of the defect’s emission frequency and its photostability by the elimination of different charge-conversion pathways.
@@
Although effective and experimentally straightforward to implement, the use of atomic and molecular adsorbates is limited by their thermal and chemical stability in the long-term Struck and D’Evelyn (1993); Thomas et al. (1992). Hence, a more robust passivation scheme is needed that retains the different benefits of using simple adsorbates, but eliminates their limitations. That this can be achieved was shown in an experimental work by Polking et al. Polking et al. (2018). The authors grew an epitaxial layer of wurtzite aluminum nitride (AlN) on a SiC surface and reported improved photostability of the positively-charged carbon antisite-vacancy defect complexes within SiC. These were promising experimental results, however, to date, there are no theoretical studies to explain how AlN passivation restores a defect’s photostability, or its effects on other SiC defects. For example, it is not known if such a passivation scheme will also improve the charge-state stability of the silicon monovacancy, which is bright only in the negatively charged state. In this first principles work, we study AlN-coverage of SiC surfaces by modeling a core-shell nanowire with a SiC core and an AlN shell. Using in SiC as a proof-of-principle defect, we demonstrate that AlN effectively removes SiC surface states from the band gap, restoring the defect’s optical properties. We also report the existence of novel silicon vacancy-based defects at the SiC-AlN interface with distinct spin and optical properties.
Computational Methods
@@
Our density functional theory (DFT)-based spin-polarized calculations were performed using the QUANTUM ESPRESSO package Giannozzi et al. (2009, 2017). We used ultrasoft pseudopotentials Vanderbilt (1990), which allow for low energy cut-offs of 50 Ry for expanding wave functions and 350 Ry for charge densities. In addition, all of the results reported in this work are for structures that were relaxed until the forces on the atoms were smaller than Ry/a.u. We used the generalized gradient potential approximation (GGA) Perdew and Yue (1986) of Perdew, Burke, and Ernzerhof (PBE) Perdew et al. (1996) to account for the exchange-correlation effects. It should be added here that we chose to use PBE functional instead of computationally expensive hybrid functionals, such as HSE06 Heyd et al. (2003, 2006). The latter functionals are known to correct for the underestimation of band gaps within PBE Cohen et al. (2008). However, in the present study, we chose PBE functional because: (a) it made our calculations for large nanostructures feasible, and (ii) in our nanostructures, the PBE-related underestimation of band gaps is compensated by the quantum confinement effects that come into play in nanostructures that result in larger band gaps. This compensation ensures a correct placement of defect states in the band gap and hence, to accurate values for quantum emission frequencies.
@@
Our structure is a core-shell NW with a 2H-SiC polytype at its core and a shell of wurtzite AlN. As done in recent studies Joshi and Dev (2022); Sayou Ngomsi et al. (2024), we chose to use the 2H-SiC polytype for the NW because it ensured that the as-created stoichiometric nanostructures were non-magnetic, with no magnetic moment on any atom. This prevented any surface-related spin from interacting with the spin of , allowing us to concentrate on the role that the finite-size and surface effects play in modifying the defect’s properties. Our calculations show the bulk lattice constants, (in-plane) and (out-of-plane) of wurtzite AlN and 2H-SiC differ by 1.3% and 2.3%, respectively. Hence, wurtzite AlN is a good choice for the passivation layer due to its small lattice-mismatch with 2H-SiC. The ideal unpassivated SiC/AlN NW, shown in Fig. 1(a), has a diameter of . The NW is periodic along the direction (c-axis), with a vacuum of about separating the images of the NW in the lateral direction (perpendicular to the c-axis), minimizing the interaction between the periodic images of the nanowire. We use a gamma-centered k-point grid of , created according to the Monkhorst-Pack scheme Monkhorst and Pack (1976). The unpassivated core-shell NW (from here on, called SiC/AlN NW) has a total of 384 atoms, with a SiC core of atoms and an AlN shell consisting of atoms. The SiC/AlN NW was further passivated with hydrogen for reasons that will be discussed in the next section. It is this passivated core-shell structure (SiC/AlN-H NW) that was used in our study. Figure 1(b) shows the cross-sectional view of the SiC/AlN-H NW with a total of 480 atoms. There are two distinct possibilities for the placement of silicon monovacancies relative to the interface between SiC and AlN: within the SiC core (i.e away from the interface) or at the interface itself. Two such sites – S1 and S2 – are shown in Figure 1(b). In this work, we used silicon vacancies at these sites to determine the efficacy of AlN in removing dangling bonds and restoring the optical properties of the monovacancies. Since the hydrogen-passivated SiC NW used in the older work by Sayou et al. Sayou Ngomsi et al. (2024) had a smaller diameter, we performed additional calculations for a hydrogen-passivated SiC NW (henceforth, called SiC-H NW), consisting of 480 atoms for the sake of direct comparison with the results for defects in SiC/AlN-H NW.
@@
In order to study the optical properties of the near-surface in SiC/AlN-H NW, we obtained the zero-phonon line (ZPL) using the SCF method Gali et al. (2009), which has been successfully employed to study excitations between spatially-localized defect states of deep-level defects in different wide band gap semiconductors Dev (2020); Joshi and Dev (2022); Narayanan and Dev (2023); Sayou Ngomsi et al. (2024). In the SCF method, the occupation of the defect states is constrained to mimic the photo-excitation process as described within the Franck-Condon model shown in Fig. 1(c). Point A in the Franck-Condon picture represents the system in the electronic ground state with the ground-state equilibrium ionic positions (generalized coordinate ). The spin-preserving, vertical optical excitation, , involves emptying of a filled defect state and the simultaneous occupation of the previously empty defect state (without structural relaxation). Point C in the Franck-Condon picture is accessed by allowing for the ionic relaxation (excited-state generalized coordinate ) that follows the new electronic configuration in the excited-state. Point D corresponds to the system with excited-state ionic positions, , and ground state electronic configuration. The total energy differences for the system at points , , and yield the following: (i) vertical transition upon absorption of a photon (), (ii) purely electronic ZPL transition (), (iii) the Stokes shift () and (iv) the vertical de-excitation upon emission of a photon ().
Results and Discussion
.1 Defect-free Core-Shell NW
@@
The spatial and energetic proximity of the surface states to the defect states allows them to interfere with the photostability of the quantum emitters Joshi and Dev (2022); Sayou Ngomsi et al. (2024). This can be avoided by passivating SiC dangling bonds within the passivation-scheme of choice, be it simple adsorbates or a coverage with another semiconductor. In principle, with the AlN-passivation of SiC one should obtain an electronic structure where the surface-states from the AlN shell are far removed from defect states (at least spatially). However, as can be seen in the density of states (DOS) plot in Fig 2(a) for the bare/unpassivated SiC/AlN NW, we have surface states in the band gap. Figures 2(b) and (c) plot the DOS projected onto the and orbitals of atoms belonging to the AlN shell and the interfacial Si and C atoms of the SiC core. A closer look at the two figures shows that it is the threefold-coordinated Al and N atoms, along with the C atoms at the interface that contribute the most to the states at the band edges of the core-shell NW. One expects to find the surface states originating from threefold-coordinated Al and N atoms at the band edges Sayou Ngomsi et al. (2024). However, the reasons for the presence of the states from the interfacial C atoms at the band edges are not self-evident. For this, we need to consider the skin-bond contraction effect, which is found in very small nanostructures Huang et al. (2008). The skin contraction occurs due to the strengthening of the bonds at the undercoordinated surface sites, which aids in a reduction of the surface energy. The outcome of the bond-contraction is evident in Fig. 1(a), which shows a marked change of hybridization from the ideal to a mixed - type for the surface atoms [compare with Fig. 1(b)]. These structural changes at the surface propagate to the interior of the NW Joshi and Dev (2022); Sayou Ngomsi et al. (2024). Since we have a thinner shell, these structural changes at the AlN surface propagate to the SiC-AlN interface, affecting the structure of SiC at the interface. The presence of surface states and the states from the interfacial atoms, in turn, reduces the band gap of the SiC/AlN NW to 2.17 eV, which is smaller than the theoretical band gap (2.31 eV) of bulk 2H-SiC. The calculated electronic structure of the SiC/AlN NW are an artifact of our core-shell model, in which the AlN shell is extremely thin. However, in order to demonstrate that AlN grown on SiC is an effective passivation scheme for improving optical properties of near-surface defects, it should be shown that the hybridization between the detrimental surface states and the defect states from in SiC is minimal. This can be achieved only if the defect states are energetically and/or spatially far removed from surface states. In other words, we need to either create a structure with much thicker shell or a structure that mimics the behavior of a thicker shell. Due to the prohibitive cost of calculations involving thicker shells, we decided to emulate a much thicker AlN shell by passivating the structure with hydrogen. A full hydrogen coverage of all threefold coordinated surface atoms (totaling 96 atoms) removed the detrimental surface and interfacial states from the band edges as seen in Fig. 2(d). As a consequence, the band gap of the resulting SiC/AlN-H NW is 2.85 eV. As expected, this is larger than the band gap for the bulk material due to quantum confinement effects that come into play in a nanostructure. For comparison, we also plot the DOS of the SiC-H NW, which shows a band gap of 2.77 eV.
.2 Silicon vacancy in passivated SiC/AlN-H NW
As discussed in the previous section, the hydrogenation of the surface of our extremely thin AlN-shell allowed us to emulate the behavior of experimentally-relevant thicknesses of AlN shells Polking et al. (2018). In doing so, we also remedied the excessive structural changes at the SiC-AlN interface, which are artifacts of the skin-contraction effects and not due to lattice mismatch. Our next step was to study properties of at the S1 and S2 sites in SiC/AlN-H. Before presenting our results for at these sites, it is instructive to see how the presence of the AlN shell may change the charge state stability. To do so, we calculated the defect formation energies as defined in earlier works Tanaka et al. (2003); Sayou Ngomsi et al. (2024) for at the S1 (interior) and S2 (interface) sites within the SiC/AlN-H NW. Figures 3 (a) and (b) are plots of defect formation energies vs electronic chemical potential (Fermi energy) referenced to the valence band maximum. These results are obtained under carbon-rich (i.e. silicon-poor) conditions, which would be the growth condition that favors formation of silicon vacancies. The highlighted regions reveal that the bright, negatively-charged state of is stable over a considerable range of n-doping for both S1 and S2 sites, although the favorable range for the S2-site is shifted to lower values of n-doping than that for S1. We would like to point out that our results for at S1-site [see Fig. 3(a)] are similar to those obtained for simple adsorbate-passivated SiC NW Sayou Ngomsi et al. (2024) and to those reported for monovacancies in bulk 4H-SiC within the HSE06 approximation Szász et al. (2015); Gordon et al. (2015). We also find that the charge state becomes favorable for a much larger doping range for the defect at S2, which might become an important consideration for highly n-doped SiC.
.2.1 at the S1 (interior) site in SiC/AlN-H NW

### v2

# Robust Aluminum Nitride Passivation of Silicon Carbide with Near-Surface Quantum Emitters for Quantum Computing and Sensing Applications

@@ Abstract
Silicon carbide (SiC) hosts a number of point defects that are being explored as single-photon emitters for quantum applications. Unfortunately, these quantum emitters lose their photostability when placed in proximity to the surface of the host semiconductor. In principle, a uniform passivation of the surface’s dangling bonds by simple adsorbates, such as hydrogen or mixed hydrogen/hydroxyl groups, should remove detrimental surface effects. However, the usefulness of atomic and molecular passivation schemes is limited by their lack of long-term chemical and/or thermal stability. In this first principles work, we use aluminum nitride (AlN) to passivate SiC surfaces in a core-shell nanowire model. By using a negatively charged silicon vacancy in SiC as the proof-of-principle quantum emitter, we show that AlN-passivation is effective in removing SiC surface states from the band gap and in restoring the defect’s optical properties. We also report the existence of a silicon vacancy-based defect at the SiC-AlN interface, which displays distinct spin and optical properties as compared to the other well-studied defects in SiC.
keywords:
[…]
@@ Introduction
Quantum bits (qubits) and single photon emitters are essential building blocks for quantum information science applications. Wide band gap semiconductors with fluorescent, spin-active point defects can furnish both qubits and single photon emitters for quantum technologies. This is particularly true for different bright defects in silicon carbide (SiC), which not only emit single photons in the favorable near-telecom range 1, 2, but they also possess addressable spins 3, 4, 5, 6 with long coherence times 5, 1, 7, 6, 4. The host material itself offers additional advantages, such as low cost, commercial availability, and ease of nanofabrication, all of which can enable scalable SiC-based quantum technologies.
@@
In order to enhance the radiative signal from defects for quantum computing and/or enhance their sensitivity for quantum sensing applications, different strategies are used, such as shallow implantation and/or nanofabrication of optical cavities in the host materials 8, 9, 10, 11, 12, 13, 14. A consequence of these strategies is the near-surface placement of the defect, which inadvertently modulates its properties. This has been reported in experiments 13, 15, 16, 17, 18 and explored in a few theoretical works 19, 20, 21, 22, 23. In particular, two recent theoretical works 22, 23 studied properties of the near-surface, negatively-charged silicon monovacancy () in a 2H-SiC nanowire (NW). In their first principles study of defects in an unpassivated NW, Joshi and Dev 22 demonstrated that the changes to the frequency of quantum emission from and the loss of the defect’s photostability result from the spatial and energetic proximity of the surface states to the defect states. Sayou Ngomsi et al. 23 further showed that the aforementioned detrimental surface effects can be remedied by passivation of the SiC surfaces with simple adsorbates, such as hydrogen or mixed hydrogen/hydroxyl groups. The resulting removal of the SiC surface states from the band gap leads to a near-perfect restoration of the defect’s emission frequency and its photostability by the elimination of different charge-conversion pathways.
@@
Although effective and experimentally straightforward to implement, the use of atomic and molecular adsorbates is limited by their thermal and chemical stability in the long-term 24, 25, 26. Hence, a more robust passivation scheme is needed that retains the different benefits of using simple adsorbates, but eliminates their limitations. That this can be achieved was shown in an experimental work by Polking et al. 27. The authors grew an epitaxial layer of wurtzite aluminum nitride (AlN) on a SiC surface and reported improved photostability of the positively-charged carbon antisite-vacancy defect complexes within SiC. These were promising experimental results, however, to date, there are no theoretical studies to explain how AlN passivation restores a defect’s photostability, or its effects on other SiC defects. For example, it is not known if such a passivation scheme will also improve the charge-state stability of the silicon monovacancy, which is bright only in the negatively charged state. In this first principles work, we study AlN-coverage of SiC surfaces by modeling a core-shell nanowire with a SiC core and an AlN shell. Using in SiC as a proof-of-principle defect, we demonstrate that AlN effectively removes SiC surface states from the band gap, restoring the defect’s optical properties. We also report the existence of novel silicon vacancy-based defects at the SiC-AlN interface with distinct spin and optical properties.
Computational Details
@@
Our density functional theory (DFT)-based spin-polarized calculations were performed using the QUANTUM ESPRESSO package 28, 29. We used ultrasoft pseudopotentials 30, which allow for low energy cut-offs of 50 Ry for expanding wave functions and 400 Ry for charge densities. In addition, all of the results reported in this work are for structures that were relaxed until the forces on the atoms were smaller than Ry/a.u. We used the generalized gradient potential approximation (GGA) 31 of Perdew, Burke, and Ernzerhof (PBE) 32 to account for the exchange-correlation effects. The PBE functional was chosen instead of the prohibitively expensive Heyd-Scuseria-Ernzerhof (HSE) hybrid functional 33, 34 because: (i) it made our calculations for large nanostructures feasible, and (ii) our emphasis is on properties that are expected to remain the same whether PBE or the computationally expensive hybrid functionals are used. For example, in this work, we explored if AlN can effectively passivate SiC surfaces, and if this passivation scheme changes the optical properties of the monovacancies. For the latter, we considered changes between the quantum emission frequencies of the monovacancies in the core-shell NW and bulk. Since such a change in emission frequency is a difference of differences in total energies, any error in our predictions was minimized.
Our structure is a core-shell NW with a 2H-SiC polytype at its core and a shell of wurtzite AlN. As done in recent studies 22, 23, we chose to use the 2H-SiC polytype for the NW because it ensured that the as-created stoichiometric nanostructures were non-magnetic, with no magnetic moment on any atom. This prevented any surface-related spin from interacting with the spin of , allowing us to concentrate on the role that the finite-size and surface effects play in modifying the defect’s properties. Our calculations show the bulk lattice constants, (in-plane) and (out-of-plane) of wurtzite AlN and 2H-SiC differ by 1.3% and 2.3%, respectively. Hence, wurtzite AlN is a good choice for the passivation layer due to its small lattice-mismatch with 2H-SiC. Note that other hexagonal polytypes, such as 4H- and 6H-SiC, are the more commonly explored hosts of quantum emitters 1, 2, 3, 4, 5, 6, 7, 35, 36, 37, 38. These hexagonal polytypes differ from each other and from 2H-SiC in the stacking of the tetrahedrally-bonded silicon and carbon atoms. The differences in stacking sequences result in different local crystal fields with distinct h- and k-sites, as well as different band gaps of the hexagonal polytypes. In turn, these differences result in variations in the excited state properties, such as the quantum emission frequencies, of their quantum emitters 3. Nevertheless, qualitatively, the results from our work on -based defects in 2H-SiC are general and should apply to other SiC polytypes as well. It should also be added that the feasibility and effectiveness of using AlN as a passivation layer for a different polytype (4H-SiC) has already been established by Polking et al. 27.
@@
The ideal unpassivated SiC/AlN NW, shown in Figure 1(a), has a diameter of Å. The NW is periodic along the direction (c-axis), with a vacuum of about Å separating the images of the NW in the lateral direction (perpendicular to the c-axis), minimizing the interaction between the periodic images of the nanowire. We use a gamma-centered k-point grid of , created according to the Monkhorst-Pack scheme 39. The unpassivated core-shell NW (from here on, called SiC/AlN NW) has a total of 384 atoms, with a SiC core of atoms and an AlN shell consisting of atoms. The surface of the as-created SiC/AlN NW has threefold- and four-fold coordinated atoms 22. The undercoordinated atoms on the surface of this NW were further passivated with hydrogen (H) atoms for reasons that will be discussed in the next section. It is this passivated core-shell structure (SiC/AlN-H NW) that was used in our study. Figure 1(b) shows the cross-sectional view of the SiC/AlN-H NW with a total of 480 atoms. Further structural details are provided in Supplementary Figures S1(a) and (b) in the Supporting Information (SI). There are two distinct possibilities for the placement of silicon monovacancies relative to the interface between SiC and AlN: within the SiC core (i.e away from the interface) or at the interface itself. Two such sites – S1 and S2 – are shown in Figure 1(b). In this work, we used silicon vacancies at these sites to determine the efficacy of AlN in removing dangling bonds and restoring the optical properties of the monovacancies. Since the hydrogen-passivated SiC NW used in the older work by Sayou et al. 23 had a smaller diameter, we performed additional calculations for a hydrogen-passivated SiC NW (henceforth called SiC-H NW), consisting of 480 atoms for sake of direct comparison with the results for defects in the SiC/AlN-H NW. Supplementary Figures S1(c) shows the cross-sectional (top) view of the SiC-H NW.
@@
In order to study the optical properties of the near-surface in SiC/AlN-H NW, we obtained the zero-phonon line (ZPL) using the SCF method 40, which has been successfully employed to study excitations between spatially-localized defect states of deep-level defects in different wide band gap semiconductors 41, 22, 42, 23. In the SCF method, the occupation of the defect states is constrained to mimic the photo-excitation process as described within the Franck-Condon model shown in Figure 1(c). Point A in the Franck-Condon picture represents the system in the electronic ground state with the ground-state equilibrium ionic positions (generalized coordinate ). The spin-preserving, vertical optical excitation, , involves emptying of a filled defect state and the simultaneous occupation of the previously empty defect state (without structural relaxation). Point C in the Franck-Condon picture is accessed by allowing for the ionic relaxation (excited-state generalized coordinate ) that follows the new electronic configuration in the excited-state. Point D corresponds to the system with excited-state ionic positions, , and ground state electronic configuration. The total energy differences for the system at points , , and yield the following: (i) vertical transition upon absorption of a photon (), (ii) purely electronic ZPL transition (), (iii) the Stokes shift () and (iv) the vertical de-excitation upon emission of a photon ().
1 Results and Discussion
1.1 Defect-free Core-Shell NW
@@
The spatial and energetic proximity of the surface states to the defect states allows them to interfere with the photostability of the quantum emitters 22, 23. This can be avoided by passivating SiC dangling bonds within the passivation-scheme of choice, be it simple adsorbates or a coverage with another semiconductor. In principle, with the AlN-passivation of SiC one should obtain an electronic structure where the surface-states from the AlN shell are far removed from defect states (at least spatially). However, as can be seen in the density of states (DOS) plot in Fig 2(a) for the bare/unpassivated SiC/AlN NW, we have surface states in the band gap. Figures 2(b) and (c) plot the DOS projected onto the and orbitals of atoms belonging to the AlN shell and the interfacial Si and C atoms of the SiC core. A closer look at the two figures shows that it is the threefold-coordinated Al and N atoms, along with the C atoms at the interface that contribute the most to the states at the band edges of the core-shell NW. One expects to find the surface states originating from threefold-coordinated Al and N atoms at the band edges 23. However, the reasons for the presence of the states from the interfacial C atoms at the band edges are not self-evident. For this, we need to consider the skin-bond contraction effect, which is found in very small nanostructures 43. The skin contraction occurs due to the strengthening of the bonds at the undercoordinated surface sites, which aids in a reduction of the surface energy. The outcome of the bond-contraction is evident in Figure 1(a), which shows a marked change of hybridization from the ideal to a mixed - type for the surface atoms [compare with Figure 1(b)]. This change in hybridization at the surface can also be seen in the Supplementary Figure S2(a). These structural changes at the surface propagate to the interior of the NW 22, 23. Since we have a thinner shell, these structural changes at the AlN surface propagate to the SiC-AlN interface, affecting the structure of SiC at the interface. The presence of surface states and the states from the interfacial atoms, in turn, reduces the band gap of the SiC/AlN NW to 2.17 eV, which is smaller than the theoretical band gap (2.31 eV) of bulk 2H-SiC. The calculated electronic structure of the SiC/AlN NW are an artifact of our core-shell model, in which the AlN shell is extremely thin. However, in order to demonstrate that AlN grown on SiC is an effective passivation scheme for improving optical properties of near-surface defects, it should be shown that the hybridization between the detrimental surface states and the defect states from in SiC is minimal. This can be achieved only if the defect states are energetically and/or spatially far removed from surface states. In other words, we need to either create a structure with much thicker shell or a structure that mimics the behavior of a thicker shell. Due to the prohibitive cost of calculations involving thicker shells, we decided to emulate a much thicker AlN shell by passivating the structure with hydrogen.
In order to mimic the behavior of a much thicker AlN shell, we carefully considered if the undercoordinated atoms on the surface should be passivated using real hydrogens with a valence charge of , or pseudohydrogens (PH) with fractional charges 44, 45, 46 of and for passivating the threefold coordinated Al and N atoms, respectively. We found that the properties of the core-shell NWs that influence the defect states – band gaps and the DOS around the band edges – remained the same whether we used real H or PH atoms [see Supplementary Figure S3]. We attribute this to the straddling of the smaller band gap of SiC by AlN’s band gap 47 as discussed in detail within Supplementary Note I. In addition, whether we use real H or PH atoms, the surface passivation restores the character of bonds between the atoms at the surface [see Supplementary Figure S2 (b)]. For the sake of completeness, we present our results for real H-passivation in the main text, while those for the PH-passivated core-shell structures are presented in the supporting information. A full hydrogen coverage of all threefold coordinated surface atoms (totaling 96 atoms) removed the detrimental surface and interfacial states from the band edges as seen in Figure 2(d). As a consequence, the band gap of the resulting SiC/AlN-H NW is 2.85 eV. As expected, this is larger than the band gap for the bulk material due to quantum confinement effects that come into play in a nanostructure. For comparison, we also plot the DOS of the SiC-H NW, which shows a band gap of 2.77 eV.
1.2 Silicon vacancy in passivated SiC/AlN-H NW
As discussed in the previous section, the hydrogenation of the surface of our extremely thin AlN-shell allowed us to emulate the behavior of experimentally-relevant thicknesses of AlN shells 27. In doing so, we also remedied the excessive structural changes at the SiC-AlN interface, which are artifacts of the skin-contraction effects and not due to lattice mismatch. Our next step was to study properties of at the S1 and S2 sites in SiC/AlN-H. Before presenting our results for at these sites, it is instructive to see how the presence of the AlN shell may change the charge state stability. To do so, we calculated the defect formation energies for at the S1 (interior) and S2 (interface) sites within the SiC/AlN-H NW. The Freysoldt-Neugebauer-Van de Walle (FNV) correction scheme 48 as implemented in the CoFFEE code 49 was used to correct for the spurious Coulomb interactions between the charged defect and its periodic images.
Figures 3 (a) and (b) are plots of defect formation energies vs electronic chemical potential (Fermi energy) referenced to the valence band maximum for the SiC/AlN-H NW. For comparison, we have also provided the formation energies of the defect (at the interior site) in SiC-H NW in Figure 3 (c). All results are obtained under carbon-rich (i.e. silicon-poor) conditions, which would be the growth condition that favors formation of silicon vacancies. The highlighted regions reveal that the bright, negatively-charged state () of is stable over a considerable range of n-doping within the SiC/AlN-H and SiC-H NWs. Also note that the charge transition levels [(0/1-) and (1-/2-)] for the defect at the S2-site are located lower in the band gap as compared to those for the S1-site defects in the SiC/AlN-H and SiC-H NWs [see Figs 3 (a) and (c)]. Although the formation energies for monovacancies in SiC/AlN-PH follow similar qualitative trends [see Supplementary Figure S4 and Supplementary Note II], the charge state for the defect at the S2-site is stable over a narrower range of chemical potential in SiC/AlN-PH than in SiC/AlN-H.
1.2.1 at the S1 (interior) site in SiC/AlN-H NW